# Numerical investigation of the interior geometry of semiclassical evaporating spherical charged black holes

### v1

@@ I Introduction
The mentioned analytical approximation was developed under a key assumption about the behavior of the RSET in the evolving metric: our assumption is that the asymptotic behavior found near the IH in RN is naturally generalized to the evolving backreaction geometry, in a certain manner that we specify at the end of subsection III.6 below.
Our goal in this work is to numerically explore this issue – more specifically, to check the validity of the above mentioned analytical approximation. To this end, we developed a numerical code which allows us to evolve the semiclassical Einstein equation in this model of spherically symmetric metric that evolves inside the evaporating semiclassical charged black hole. We work in double-null coordinates, using the outgoing null coordinate $u$ and the ingoing null coordinate $v$. The resulting equations are partial differential equations for the two unknown functions $R(u,v)$ and $S(u,v)$ (to be defined later), which fully define the spherically symmetric metric. We begin our numerical simulation close to the EH, and evolve the Einstein’s evolution equations towards the future, until we get deep in the near-IH region. Our numerical simulation confirms the validity of the mentioned analytical approximation.
II Theoretical background
II.1 Spherically symmetric Einstein equations
@@
$d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ (II.2)
We’ll use geometrized unit system where $c=1$ and $G=1$. In this unit system, there’s an equivalence between distance, time and mass units. This metric depends on two unknown functions: $S(u,v)$ and $r(u,v)$. $r(u,v)$ has units of length, and so has $L_{0}$. $S(u,v)$ is dimensionless. We’re free to choose $L_{0}$ as we wish. We define
$R(u,v)=r^{2}(u,v)$ (II.3)
@@
$G_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (II.4)
$G_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (II.5)
$G_{uv}=\frac{L_{0}e^{S}}{R^{3/2}}+\frac{R_{,uv}}{R}$ (II.6)
$G_{\theta\theta}=-\frac{1}{2L_{0}}e^{-S}\sqrt{R}\left(R_{,uv}+2RS_{,uv}\right)$ (II.7)
@@
$R_{,uv}=-\frac{L_{0}e^{S}}{\sqrt{R}}+8\pi RT_{uv}$ (II.9)
$S_{,uv}=\frac{L_{0}e^{S}}{2R^{\frac{3}{2}}}-8\pi\left(\frac{1}{2}T_{uv}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}\right)$ (II.10)
@@
$8\pi T_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (II.11)
$8\pi T_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (II.12)
The equations for $R_{,uu}$ and $R_{,vv}$, entailed in Eqs. (II.11,II.12), are the constraint equations, and the equations (II.9,II.10) for $R_{,uv}$$S_{,uv}$ are the evolution equations.
@@
$T_{\mu\nu}=T_{\mu\nu}^{(EM)}+T_{\mu\nu}^{(SC)}$ (II.13)
II.2 Solution to Maxwell’s equations
@@
$F_{\alpha\beta;\gamma}+F_{\beta\gamma;\alpha}+F_{\gamma\alpha;\beta}=0$ (II.14)
@@
$F_{\alpha\beta,\gamma}+F_{\beta\gamma,\alpha}+F_{\gamma\alpha,\beta}=0$ (II.15)
It is easy to see the equation is trivially satisfied. The source-free inhomogeneous equation is:
$0=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (II.16)
@@
$\left(\sqrt{|g|}F^{uv}\right)_{,v}=0$ (II.18)
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,v}=0$ (II.19)
Taking the $\alpha=v$ component of the equation Eq. (II.16) yields an analogous result:
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,u}=0$ (II.20)
From these two equations we get that $r^{2}|g_{uv}|F^{uv}$ is constant, which we call $Q$.
@@
$4\pi J^{\alpha}=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (II.25)
@@
$4\pi J^{u}=\frac{1}{r^{2}\sin\theta|g_{uv}|}\left(r^{2}\sin\theta|g_{uv}|\frac{Q(u,v)}{r^{2}|g_{uv}|}\right)_{,v}=\frac{Q_{,v}}{r^{2}|g_{uv}|}$ (II.26)
@@
$4\pi J_{u}=g_{uv}\frac{-Q_{,u}}{r^{2}|g_{uv}|}=\frac{Q_{,u}}{r^{2}}$ (II.28)
$4\pi J_{v}=-\frac{Q_{,v}}{r^{2}}$ (II.29)
II.3 The electromagnetic stress-energy tensor
@@
$T_{\mu\nu}^{(EM)}=\frac{1}{4\pi}\left(F_{\mu\lambda}F_{\nu}^{\;\lambda}-\frac{1}{4}g_{\mu\nu}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (II.30)
@@
$F_{\lambda\kappa}F^{\lambda\kappa}=2F_{uv}F^{uv}=-2\frac{Q}{r^{2}|g_{uv}|}\frac{Q}{r^{2}}|g_{uv}|=-\frac{2Q^{2}}{r^{4}}$ (II.31)
@@
$T_{uv}^{(EM)}$ $=\frac{1}{4\pi}\left(F_{uv}g^{uv}F_{vu}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (II.32)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}|g_{uv}|^{2}g^{uv}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (II.33)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}g_{uv}+\frac{1}{4}g_{uv}\frac{2Q^{2}}{r^{4}}\right)=-\frac{1}{8\pi}\frac{Q^{2}}{r^{4}}g_{uv}$ (II.34)
@@
$T_{\theta\theta}^{(EM)}=\frac{1}{4\pi}\left[-\frac{1}{4}g_{\theta\theta}\left(-\frac{2Q^{2}}{r^{4}}\right)\right]=\frac{1}{8\pi}\frac{Q^{2}}{r^{2}}$ (II.37)
Note that $T_{vv}^{(EM)}=T_{uu}^{(EM)}=0$, therefore $T_{uu}=T_{uu}^{(SC)}$ and $T_{vv}=T_{vv}^{(SC)}$. We will hereafter refer to $T_{uu}$ and $T_{vv}$ without the “(SC)(SC)” mark.
Using Einstein’s equation $G_{\mu\nu}=8\pi T_{\mu\nu}$, and assuming the only stress-energy tensor is the electromagnetic one, the evolution equations in the presence of charge are, based on equations Eq. (II.9,II.10)
$R_{,uv}=-\frac{e^{S}L_{0}}{\sqrt{R}}+e^{S}L_{0}\frac{Q^{2}}{R^{\frac{3}{2}}}=e^{S}L_{0}F_{1}(R)$ (II.38)
$S_{,uv}=\frac{e^{S}L_{0}}{2R^{\frac{3}{2}}}-\frac{3}{2}e^{S}L_{0}\frac{Q^{2}}{R^{\frac{5}{2}}}=e^{S}L_{0}F_{2}(R)$ (II.39)
@@
$F_{1}(R)=-\frac{1}{\sqrt{R}}+\frac{Q^{2}}{R^{\frac{3}{2}}}$ (II.40)
$F_{2}(R)=\frac{1}{2R^{\frac{3}{2}}}-\frac{3}{2}\frac{Q^{2}}{R^{\frac{5}{2}}}$ (II.41)
@@
$F_{2}(R)=\frac{d}{dR}F_{1}(R)$ (II.42)
In view of this property, the PDE system (II.38,II.39) belongs to the general class of PDE systems investigated in Ref. [24].
Note that we here assumed that the only stress-energy tensor is the electromagnetic one. We will later break this assumption, by taking into account the non-zero $T_{\mu\nu}^{(SC)}$. This will lead to new source terms in the evolution equations, see subsection II.8.
Taking the constraint equation Eq. (II.12) we get:
$\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$ (II.43)
@@
$\left(\widetilde{T}_{uu}\right)_{,v}=0$ (II.45)
To prove this, we differentiate Eq. (II.44) with respect to $v$, and use the evolution equations Eq. (II.9,II.10):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (II.46)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-\left[e^{S}L_{0}F_{1}(R)\right]_{,u}$ (II.47)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-e^{S}L_{0}F_{1}(R)S_{,u}-e^{S}L_{0}\frac{dF_{1}(R)}{dR}R_{,u}=0$ (II.48)
In the last equation we used the relation Eq. (II.42) between $F_{1}$ and $F_{2}$. Note that we assumed $Q$ is constant at this stage. When $Q$ changes the situation changes and we address it in section III.6.
@@
$\left(\widetilde{T}_{vv}\right)_{,u}=0$ (II.49)
Note, however, that when we add the semiclassical source terms Eqs. (II.45, II.49) will not be valid.
II.4 Reissner–Nordström metric
@@
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dt^{2}+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}$ (II.50)
@@
$r_{+}=M+\sqrt{M^{2}-Q^{2}}$ (II.51)
$r_{-}=M-\sqrt{M^{2}-Q^{2}}$ (II.52)
$\kappa_{+}=\frac{r_{+}-r_{-}}{2r_{+}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(\sqrt{M^{2}-Q^{2}}+M\right)^{2}}$ (II.53)
$\kappa_{-}$ $=\frac{r_{+}-r_{-}}{2r_{-}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(M-\sqrt{M^{2}-Q^{2}}\right)^{2}}$ (II.54)
$=\frac{1}{M}\frac{\sqrt{1-\left(\frac{Q}{M}\right)^{2}}}{\left(1-\sqrt{1-\left(\frac{Q}{M}\right)^{2}}\right)^{2}}$ (II.55)
@@
$f\equiv\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}$ (II.56)
@@
II.5 Ingoing Eddington coordinates

### v2

@@ 1 Introduction
The mentioned analytical approximation was developed under a key assumption about the behavior of the RSET in the evolving metric: our assumption is that the asymptotic behavior found near the IH in RN is naturally generalized to the evolving backreaction geometry, in a certain manner that we specify at the end of subsection 3.6 below.
Our main goal in this work is to numerically explore this issue – more specifically, to check the validity of the above mentioned analytical approximation. To this end, we developed a numerical code (published as [23]) which allows us to evolve the semiclassical Einstein equation in this model of spherically symmetric metric that evolves inside the evaporating semiclassical charged black hole. We work in double-null coordinates, using the outgoing null coordinate $u$ and the ingoing null coordinate $v$. The resulting equations are partial differential equations for the two unknown functions $R(u,v)$ and $S(u,v)$ (to be defined later), which fully define the evolving spherically symmetric metric. We begin our numerical simulation close to the EH, and evolve the Einstein’s evolution equations towards the future, until we get deep into the near-IH region. The numerical simulation assumes a specific form for the mass $M(v)$ of the evaporating black hole as a function of the ingoing Eddington coordinate $v$, and that the charge-to-mass ratio $\frac{Q}{M}$ at the EH remains constant upon evaporation – as well as a form for the source terms which incorporate the semiclassical contributions to $T_{uv}$ and $T_{\theta\theta}$, and for the current $J_{u}$  [^2] In practice, we choose a function $Q(u,v)$ which determines the current components $J_{u}$ and $J_{v}$. Our numerical simulation confirms the validity of the mentioned analytical approximation.
Part I Theoretical overview
2 Theoretical background
2.1 Spherically symmetric Einstein equations
@@
$d\Omega^{2}=d\theta^{2}+\sin^{2}\theta d\phi^{2}$ (2.2)
We’ll use geometrized unit system where $c=1$ and $G=1$. In this unit system, there is an equivalence between distance, time and mass units. This metric depends on two unknown functions: $S(u,v)$ and $r(u,v)$. $r(u,v)$ has units of length, and so has $L_{0}$. $S(u,v)$ is dimensionless. The coordinates $u$,$v$ have dimensions of length. We’re free to choose $L_{0}$ as we wish. We define
$R(u,v)=r^{2}(u,v)$ (2.3)
@@
$G_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (2.4)
$G_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (2.5)
$G_{uv}=\frac{L_{0}e^{S}}{R^{3/2}}+\frac{R_{,uv}}{R}$ (2.6)
$G_{\theta\theta}=-\frac{1}{2L_{0}}e^{-S}\sqrt{R}\left(R_{,uv}+2RS_{,uv}\right)$ (2.7)
@@
$R_{,uv}=-\frac{L_{0}e^{S}}{\sqrt{R}}+8\pi RT_{uv}$ (2.9)
$S_{,uv}=\frac{L_{0}e^{S}}{2R^{\frac{3}{2}}}-8\pi\left(\frac{1}{2}T_{uv}+\frac{L_{0}e^{S}}{R^{\frac{3}{2}}}T_{\theta\theta}\right)$ (2.10)
@@
$8\pi T_{uu}=\frac{1}{R}\left(R_{,u}S_{,u}-R_{,uu}\right)$ (2.11)
$8\pi T_{vv}=\frac{1}{R}\left(R_{,v}S_{,v}-R_{,vv}\right)$ (2.12)
The equations for $R_{,uu}$ and $R_{,vv}$, entailed in Eqs. (2.11,2.12), are the constraint equations, and the equations (2.9,2.10) for $R_{,uv}$$S_{,uv}$ are the evolution equations.
@@
$T_{\mu\nu}=T_{\mu\nu}^{(EM)}+T_{\mu\nu}^{(SC)}$ (2.13)
2.2 Solution to Maxwell’s equations
@@
$F_{\alpha\beta;\gamma}+F_{\beta\gamma;\alpha}+F_{\gamma\alpha;\beta}=0$ (2.14)
@@
$F_{\alpha\beta,\gamma}+F_{\beta\gamma,\alpha}+F_{\gamma\alpha,\beta}=0$ (2.15)
It is easy to see the equation is trivially satisfied.
The source-free inhomogeneous equation is:
$0=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (2.16)
@@
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,v}=0$ (2.18)
Taking the $\alpha=v$ component of the equation Eq. (2.16) yields an analogous result:
$\left(r^{2}|g_{uv}|F^{uv}\right)_{,u}=0$ (2.19)
From these two equations we get that $r^{2}|g_{uv}|F^{uv}$ is constant, which we call $Q$. $Q$ is the electric charge.
@@
$4\pi J^{\alpha}=F_{\quad;\beta}^{\alpha\beta}=\frac{1}{\sqrt{|g|}}\left(\sqrt{|g|}F^{\alpha\beta}\right)_{,\beta}$ (2.24)
@@
$4\pi J^{u}=\frac{1}{r^{2}\sin\theta|g_{uv}|}\left(r^{2}\sin\theta|g_{uv}|\frac{Q(u,v)}{r^{2}|g_{uv}|}\right)_{,v}=\frac{Q_{,v}}{r^{2}|g_{uv}|}$ (2.25)
@@
$4\pi J_{u}=g_{uv}\frac{-Q_{,u}}{r^{2}|g_{uv}|}=\frac{Q_{,u}}{r^{2}}$ (2.27)
$4\pi J_{v}=-\frac{Q_{,v}}{r^{2}}$ (2.28)
2.3 The electromagnetic stress-energy tensor
@@
$T_{\mu\nu}^{(EM)}=\frac{1}{4\pi}\left(F_{\mu\lambda}F_{\nu}^{\;\lambda}-\frac{1}{4}g_{\mu\nu}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.29)
@@
$F_{\lambda\kappa}F^{\lambda\kappa}=2F_{uv}F^{uv}=-2\frac{Q}{r^{2}|g_{uv}|}\frac{Q}{r^{2}}|g_{uv}|=-\frac{2Q^{2}}{r^{4}}$ (2.30)
@@
$T_{uv}^{(EM)}$ $=\frac{1}{4\pi}\left(F_{uv}g^{uv}F_{vu}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.31)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}|g_{uv}|^{2}g^{uv}-\frac{1}{4}g_{uv}F_{\lambda\kappa}F^{\lambda\kappa}\right)$ (2.32)
$=\frac{1}{4\pi}\left(-\frac{Q^{2}}{r^{4}}g_{uv}+\frac{1}{4}g_{uv}\frac{2Q^{2}}{r^{4}}\right)=-\frac{1}{8\pi}\frac{Q^{2}}{r^{4}}g_{uv}$ (2.33)
@@
$T_{\theta\theta}^{(EM)}=\frac{1}{4\pi}\left[-\frac{1}{4}g_{\theta\theta}\left(-\frac{2Q^{2}}{r^{4}}\right)\right]=\frac{1}{8\pi}\frac{Q^{2}}{r^{2}}$ (2.36)
Because of spherical symmetry, $T_{\phi\phi}=T_{\theta\theta}\sin^{2}\theta$, for both the electromagnetic stress-energy tensor and the semi-classical stress energy tensor. Note that $T_{vv}^{(EM)}=T_{uu}^{(EM)}=0$, therefore $T_{uu}=T_{uu}^{(SC)}$ and $T_{vv}=T_{vv}^{(SC)}$. We will hereafter refer to $T_{uu}$ and $T_{vv}$ without the “(SC)(SC)” mark.
Using Einstein’s equation $G_{\mu\nu}=8\pi T_{\mu\nu}$, and assuming the only stress-energy tensor is the electromagnetic one, the evolution equations in the presence of charge are, based on equations Eq. (2.9,2.10)
$R_{,uv}=-\frac{e^{S}L_{0}}{\sqrt{R}}+e^{S}L_{0}\frac{Q^{2}}{R^{\frac{3}{2}}}=e^{S}L_{0}F_{1}(R)$ (2.37)
$S_{,uv}=\frac{e^{S}L_{0}}{2R^{\frac{3}{2}}}-\frac{3}{2}e^{S}L_{0}\frac{Q^{2}}{R^{\frac{5}{2}}}=e^{S}L_{0}F_{2}(R)$ (2.38)
@@
$F_{1}(R)=-\frac{1}{\sqrt{R}}+\frac{Q^{2}}{R^{\frac{3}{2}}}$ (2.39)
$F_{2}(R)=\frac{1}{2R^{\frac{3}{2}}}-\frac{3}{2}\frac{Q^{2}}{R^{\frac{5}{2}}}$ (2.40)
@@
$F_{2}(R)=\frac{d}{dR}F_{1}(R)$ (2.41)
In view of this property, the PDE system (2.37,2.38) belongs to the general class of PDE systems investigated in Ref. [24].
Note that we here assumed that the only stress-energy tensor is the electromagnetic one. We will later break this assumption, by taking into account the non-zero $T_{\mu\nu}^{(SC)}$. This will lead to new source terms in the evolution equations, see subsection 2.8.
Taking the constraint equation Eq. (2.12) we get:
$\widetilde{T}_{vv}\equiv 8\pi r^{2}T_{vv}=R_{,v}S_{,v}-R_{,vv}$ (2.42)
@@
$\left(\widetilde{T}_{uu}\right)_{,v}=0$ (2.44)
To prove this, we differentiate Eq. (2.43) with respect to $v$, and use the evolution equations Eq. (2.9,2.10):
$\left(\widetilde{T}_{uu}\right)_{,v}$ $=R_{,uv}S_{,u}+R_{,u}S_{,uv}-\left(R_{,uv}\right)_{,u}$ (2.45)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-\left[e^{S}L_{0}F_{1}(R)\right]_{,u}$ (2.46)
$=e^{S}L_{0}F_{1}(R)S_{,u}+e^{S}L_{0}F_{2}(R)R_{,u}-e^{S}L_{0}F_{1}(R)S_{,u}-e^{S}L_{0}\frac{dF_{1}(R)}{dR}R_{,u}=0$ (2.47)
In the last equation we used the relation Eq. (2.41) between $F_{1}$ and $F_{2}$. Note that we assumed $Q$ is constant at this stage. When $Q$ changes the situation changes and we address it in section 3.6.
@@
$\left(\widetilde{T}_{vv}\right)_{,u}=0$ (2.48)
Note, however, that when we add the semiclassical source terms Eqs. (2.44, 2.48) will not be valid.
2.4 Reissner–Nordström metric
@@
$ds^{2}=-\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)dt^{2}+\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)^{-1}dr^{2}+r^{2}d\Omega^{2}$ (2.49)
@@
$r_{+}=M+\sqrt{M^{2}-Q^{2}}$ (2.50)
$r_{-}=M-\sqrt{M^{2}-Q^{2}}$ (2.51)
$\kappa_{+}=\frac{r_{+}-r_{-}}{2r_{+}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(\sqrt{M^{2}-Q^{2}}+M\right)^{2}}$ (2.52)
$\kappa_{-}$ $=\frac{r_{+}-r_{-}}{2r_{-}^{2}}=\frac{\sqrt{M^{2}-Q^{2}}}{\left(M-\sqrt{M^{2}-Q^{2}}\right)^{2}}$ (2.53)
$=\frac{1}{M}\frac{\sqrt{1-\left(\frac{Q}{M}\right)^{2}}}{\left(1-\sqrt{1-\left(\frac{Q}{M}\right)^{2}}\right)^{2}}$ (2.54)
@@
$f\equiv\left(1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\right)=\frac{\left(r-r_{-}\right)\left(r-r_{+}\right)}{r^{2}}$ (2.55)
@@
2.5 Ingoing Eddington coordinates